OpenStudy (anonymous):

NEED HELP! find the largest possible number b for which|x-3| < b implies |9-3x|<0.1

OpenStudy (anonymous):

@Solongos Can you convert the |9-3x| into |x-3| ?

OpenStudy (anonymous):

times by 3?

OpenStudy (anonymous):

no i cant

OpenStudy (anonymous):

You're close! pay attention to the sign!

OpenStudy (anonymous):

it could be 9+3x or -9-3x right? then divide them by 3

OpenStudy (anonymous):

|9-3x| = -3 |x-3|

OpenStudy (anonymous):

ohh i wrote the question wrong 1sec

OpenStudy (anonymous):

no its right

OpenStudy (anonymous):

Given |9-3x| < 0.1 -> -3 |x-3| < .1 Agree?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

oh then it will be -.1/3 ?

OpenStudy (anonymous):

=> |x-3| > - .1/3 Ok so far?

OpenStudy (anonymous):

yes ok so far

OpenStudy (anonymous):

no i don't know

OpenStudy (anonymous):

making it two different equation ?

OpenStudy (anonymous):

3+x<-.1/3 x < -4-.1/3 and -3-x<-.1/3 x>-3+.1/3

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

x>-.1/3+3 that is x>2.96667 and x<.1/3+3 that is x<3.03333

OpenStudy (anonymous):

they both says they are wrong on my web homework

OpenStudy (anonymous):

largest possible number would be 3.03333 right?

OpenStudy (anonymous):

.1/3+3= .1/3+9/3=9.1/3 right? and that is 3.03333333333333

OpenStudy (anonymous):

it says that is not the right answer on my we homework. we must have did something wrong

OpenStudy (anonymous):

Hold on, you don't have to get rid of absolute value :)

OpenStudy (anonymous):

3b?

OpenStudy (anonymous):

3.1!!!

OpenStudy (anonymous):

.1/-3!!!

OpenStudy (anonymous):

perfect it was very simple

OpenStudy (anonymous):

that was the right answer!! thank you

OpenStudy (anonymous):

i have one last question can you help me?

OpenStudy (anonymous):

I'm glad it's simple :)

OpenStudy (anonymous):

Just open the new post per question :)

OpenStudy (anonymous):

ok